OpenStudy (anonymous):

-8 – 52 ÷ 13 x 3^2 + 5 what are the steps?

OpenStudy (anonymous):

try to work through it in the same way as the last one on your own first

OpenStudy (anonymous):

Go back to the order of operations and try it one step at a time

OpenStudy (anonymous):

52 doesn't go into 117 evenly im stuck at that point..

OpenStudy (anonymous):

ah, ok

OpenStudy (anonymous):

-8 – 52 ÷ 13 x 3^2 + 5 3 * 3 = 9 9 * 13 = 117 52 / 117 = 4

OpenStudy (anonymous):

this time the answer won't be a whole number, but you're on the right track. -8 + 52/117 + 5 is a good place to be

OpenStudy (anonymous):

we can add -8 and 5 to simplify it to -3+52/117

OpenStudy (anonymous):

which comes out to -23/9 as a simplified fraction

OpenStudy (anonymous):

So what are the steps after 9 * 13 = 117 further going how I have it set up?

OpenStudy (anonymous):

-8 – 52 ÷ 13 x 3^2 + 5 3 * 3 = 9 9 * 13 = 117 52 / 117 = 2.25 2.25 + 5 = 7.25 -8 – 7.25 = -15.25 I think this is wrong?

OpenStudy (anonymous):

do 52÷13 before x9 Division and Multiplication are done whichever is on the left first. Final answer = -39